# Convolution-weighting method for the physics-informed neural network: A Primal-Dual Optimization Perspective

### v1

@@ 1 Introduction
Physics-Informed Neural Networks (PINNs) [1, 2] have established themselves as a universal framework for integrating physical laws with neural networks. By encoding governing Partial Differential Equations (PDEs) into neural network loss functions through automatic differentiation, PINNs enable solving both forward and inverse problems without requiring dense observational data. This paradigm leverages two synergistic components: the expressivity of neural networks to approximate complex solution spaces, and the regularization provided by PDE residuals and boundary/initial conditions. Its effectiveness has been demonstrated across diverse domains, from heat transfer [3, 4, 5] and solid mechanics [6, 7] to stochastic systems [8, 9] and uncertainty quantification [10, 11, 12].
@@
Recent advances in PINN training strategies primarily address loss imbalance through two complementary approaches. The first involves adaptive spatiotemporal resampling of collocation points, where regions exhibiting sharp gradients or high residuals—such as shock interfaces [22, 23] or turbulent boundary layers [24, 25]—are progressively allocated denser sampling points. For example, Lu et al. [26] developed a threshold-dependent method for the selection of new training points, whereas Wu et al. [27] devised a probability density function originated from residuals to enhance sampling efficiency. Moreover, building upon Wu et al.’s methodology, Gao et al. [28] define a “failure region” within PINN and perform resampling within this region.

### v2

@@ 1 Introduction
Physics-Informed Neural Networks (PINNs) [1, 2] have established themselves as a universal framework for integrating physical laws with neural networks. By encoding governing Partial Differential Equations (PDEs) into neural network loss functions through automatic differentiation, PINNs enable solving both forward and inverse problems without requiring dense observational data. This paradigm leverages two synergistic components: the expressivity of neural networks to approximate complex solution spaces, and the regularization provided by PDE residuals and boundary/initial conditions. Its effectiveness has been demonstrated across diverse domains, from heat transfer [3, 4, 5] and solid mechanics [6, 7] to stochastic systems [8, 9] and uncertainty quantification [10, 11, 12, 13, 14].
@@
Recent advances in PINN training strategies primarily address loss imbalance through two complementary approaches. The first involves adaptive spatiotemporal resampling of collocation points, where regions exhibiting sharp gradients or high residuals—such as shock interfaces [24, 25] or turbulent boundary layers [26, 27]—are progressively allocated denser sampling points. For example, Lu et al. [28] developed a threshold-dependent method for selecting new training points, whereas Wu et al. [29] devised a probability density function derived from residuals to enhance sampling efficiency. Moreover, building upon Wu et al.’s methodology, Gao et al. [30] define a “failure region” within PINN and perform resampling within this region. Gao et al. [31] further extend this line of inquiry by employing active learning for resampling, with a particular focus on high-dimensional cases.
@@ 4.3 Viscous Burgers equation
4.4 Unsteady Cylinder Flow
The Navier–Stokes (NS) equations stand as cornerstone frameworks for tackling an extensive array of real-world challenges rooted in the study of fluid motion. In this section, we focus on the incompressible flows passing through a cylinder, which is a unsteady 2D NS equation as follows:
$\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}$ $=-\frac{1}{\rho}\nabla p+\nu\nabla^{2}\mathbf{u},$ (32)
$\nabla\cdot\mathbf{u}$ $=0,$ (33)
where $\mathbf{u}=(u,v)$ represents the time-dependent velocity field. We choose the fluid density $\rho=1$ and $\nu=\dfrac{1}{Re}$, targeting a high Reynolds number $Re=3,900$. The Reynolds number, defined as $Re\equiv\frac{U_{\infty}D}{\nu}=3900$, with free-stream velocity ${U_{\infty}}=1$ and cylinder diameter $D=1$, characterizes flow turbulence and complexity. In this case, the flow exhibits strongly unsteady, chaotic behavior with irregular vortex shedding and transitional turbulence—a regime notoriously challenging for PINNs.
We follow the same experimental settings from the public repository [43]: The computational spatial domain is set to be $(x.y)\in\Omega=[1,5]\times[-2,2]$ and the temporal domain $t\in[0,42.93]$. The cylinder is centered at $(0,0)$ with diameter $D=1$. Rather than relying on boundary or initial condition data, often exploited in PINNs to simplify training, we use sparse in-domain measurements: 36 sensors collect 100 snapshots each, yielding 3,600 labeled data points for the data loss term. This setup mimics real-world scenarios where boundary data may be inaccessible, testing the model’s ability to generalize from limited internal observations. The sensor placement is visualized in Fig. 11.
For the physics-informed loss, we set $N_{f}=10,000$ collocation points to enforce the NS equations. The global weights for data loss and residual loss are $\lambda_{\text{data}}=10$ and $\lambda_{F}=1$ to balance the contribution of the two loss terms. The network architecture consists of 10 hidden layers, each with 32 neurons. We train the model using the Adam optimizer with an initial learning rate of $0.001$, paired with an exponential learning rate scheduler. The scheduler applies a decay factor of $0.999$ every $100$ iterations.
To evaluate performance, we randomly select 90,000 validation points from the full computational domain (not limited to sensor locations) provided in [43], ensuring we assess generalization beyond training data. As shown in Fig. 12, CWP allows PINN to achieve accurate predictions for the high-Reynolds-number NS equations at the final time slice ($t=42.93$s).
We further experiment with the other three algorithms under the same setting. The numerical results for the comparison are listed in Table 4. Consistently, CWP outperforms the other methods across all metrics. For the velocity components ($u$ and $v$), CWP reduces errors by nearly $50\%$ compared to RBA, SA, and LA. Notably, pressure $p$ errors remain higher than velocity errors across all methods, a known challenge in PINNs due to pressure’s indirect coupling to momentum in incompressible flows [1, 44, 45]. However, CWP still achieves a significant reduction in pressure error relative to the other three methods. These comparisons highlight the effectiveness of our approach in addressing the unique challenges of high-Reynolds-number flows.
Figure 11: Location of sensors (marked with black crosses) in the computational domain ($\Omega=[1,5]\times[-2,2]$). The gray circle represents a cylinder with a diameter of $D=1$ and is centered at the origin. These sensors capture observed data across 100 snapshots, which are used for training the model.
Table 4: Relative $L^{2}$ error of various algorithms on the 2D unsteady NS equation after 20,000 training iterations.
CWP RBA SA LA
x-velocity $u$ $2.06\times 10^{-2}$ $4.31\times 10^{-2}$ $5.84\times 10^{-2}$ $5.01\times 10^{-2}$
y-velocity $v$ $5.64\times 10^{-2}$ $7.10\times 10^{-2}$ $8.31\times 10^{-2}$ $7.66\times 10^{-2}$
pressure $p$ $8.15\times 10^{-2}$ $1.15\times 10^{-1}$ $1.01\times 10^{-1}$ $1.07\times 10^{-1}$
velocity $\sqrt{u^{2}+v^{2}}$ $2.29\times 10^{-2}$ $5.25\times 10^{-2}$ $5.74\times 10^{-2}$ $4.93\times 10^{-2}$
Figure 12: Heatmaps at the final time slice $t=42.93$s of the exact solution (left column), predicted solution (middle column, obtained via CWP), and point-wise error (right column) for $u(x,y)$ (top row), $v(x,y)$ (middle row), and $p(x,y)$ (bottom row) for the unsteady NS equation.